OpenStudy (anonymous):

multiply (z-x)(z^2+zx+x^2) How do I do this ?

OpenStudy (anonymous):

\[(z-x)(z^2+zx+x^2) \]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Multiply each of the terms of the binomial, (z-x), by each of the terms of the trinomial, (^2+zx+x^2).

OpenStudy (anonymous):

Or you could distribute each of the terms of the binomial like this:z(z^2+zx+x^2)+-x(z^2+zx+x^2). Then you just multiply and add/subtract like terms.

OpenStudy (anonymous):

ooooh.. thanks !

OpenStudy (anonymous):

mmm what is z*2x?

OpenStudy (anonymous):

2xz

OpenStudy (anonymous):

\[(a+b) (f+g+h)=a f+a g+a h+b f+b g+b h \]

OpenStudy (xxkesha):

\[(z-x) (z^2 +zx+x^2)=z^3+ z^2x+zx^2-z^2x-zx^2-z^3= z^3-x^3\]

candycove (candycove):

Are you my tooth fairy?

OpenStudy (anonymous):

ahahah maybe

candycove (candycove):

:D

OpenStudy (anonymous):

((: